### v1

# Variational Autoencoders for Efficient Simulation-Based Inference

Abstract
We present a generative modeling approach based on the variational inference framework for likelihood-free simulation-based inference. The method leverages latent variables within variational autoencoders to efficiently estimate complex posterior distributions arising from stochastic simulations. We explore two variations of this approach distinguished by their treatment of the prior distribution. The first model adapts the prior based on observed data using a multivariate prior network, enhancing generalization across various posterior queries. In contrast, the second model utilizes a standard Gaussian prior, offering simplicity while still effectively capturing complex posterior distributions. We demonstrate the efficacy of these models on well-established benchmark problems, achieving results comparable to flow-based approaches while maintaining computational efficiency and scalability.
1 Introduction
The simulation-based inference (SBI) problem involves statistical inference of parameters $\boldsymbol{\theta}$ of a stochastic simulation model $M(\boldsymbol{\theta},\xi)$ containing random states $\xi$, from observed data $\boldsymbol{y}$ (Cranmer et al., 2020). Once calibrated, the simulation model can be used to reason about, analyze and explain the observed data $\boldsymbol{y}$ in the context of the corresponding physical process (Lavin et al., 2021). Several challenges arise in SBI owing to model stochasticity and potential multi-valuedness, where different sets of parameter values can produce similar observations, or similar parameters may lead to varied outputs. Consequently, the task is to infer the posterior distribution $p(\boldsymbol{\theta}\mid\mathbf{y})$ in a Bayesian setting.
A user-defined prior distribution $p(\boldsymbol{\theta})$ encodes domain knowledge or assumptions about plausible parameter values. The likelihood function $p(\mathbf{y}\mid\boldsymbol{\theta})$ is however, often intractable due to intricate stochastic dynamics, high-dimensional integration, or latent variables making analytical computation impractical (Sisson et al., 2018). Although a closed-form expression for the likelihood function $p(\mathbf{y}\mid\boldsymbol{\theta})$ is unavailable, samples from this distribution can be generated by evaluating the simulator $M$ with varying $\xi$ for different $\boldsymbol{\theta}$ drawn from the prior, effectively yielding samples from the joint distribution $(\boldsymbol{\theta},\mathbf{y})\sim p(\boldsymbol{\theta},\mathbf{y})$.
Deep learning based approaches that directly approximate the model likelihood, or the posterior distribution have gained increasing popularity in the SBI setting (Zammit-Mangion et al., 2024). Several approaches involve flow-based models, which are potentially compute-intensive and also restrict the design space of neural networks to invertible functions with computationally efficient Jacobian calculations (Rezende and Mohamed, 2015). This constraint can limit the flexibility of the model in handling more complex data structures. Alternative approaches like Generative Adversarial Training for SBI (GATSBI) (Ramesh et al., 2022) use an adversarial network, where a generator produces simulations that mimic observed data, and a discriminator distinguishes between real and simulated data. However, the adversarial approach can suffer from training instability, mode collapse, and difficulties in balancing the generator and discriminator networks (Arjovsky and Bottou, 2022; Arjovsky et al., 2017), complicating training and limiting robustness.
Contribution: We propose a variational inference approach for simulation-based inference, utilizing a Conditional Variational Autoencoder (C-VAE) architecture (Kingma and Welling, 2014; Sohn et al., 2015) that is simple, computational efficient, scalable, interpretable and robust. As VAEs are not constrained by the invertibility requirement, they allow seamless integration of various neural network architectures without significant modifications as opposed to flow-based models. VAEs also avoid the instability associated with GANs due to adversarial learning dynamics. As a result, C-VAEs can handle complex data structures and dependencies more effectively, making them suitable for a wider range of tasks without requiring external summary networks (Cranmer et al., 2020).
The paper is organized as follows. Section 2 outlines the related work. In Section 3, we provide a formal problem definition, followed by a detailed explanation of the proposed architecture in Sections 3.1 and 3.2. A comprehensive comparison on benchmark test problems is presented in Section 4 and in the Appendix. Section 5 discusses the results, while Section 6 concludes the paper by summarizing the key findings and further advancements.
2 Related Work
A vast majority of SBI approaches based on deep learning aim to either approximate the likelihood (Hermans et al., 2020), or focus on directly approximating the posterior (Papamakarios and Murray, 2016; Lueckmann et al., 2017; Greenberg et al., 2019). Additionally, there exist methods that target both the likelihood and posterior simultaneously (Radev et al., 2023a; Glöckler et al., 2022; Wiqvist et al., 2021)
Neural Posterior Estimation (NPE) (Papamakarios and Murray, 2016) was one of the earliest approaches towards one-step posterior estimation, introducing the use of conditional normalizing flows to approximate the posterior distribution. By incorporating techniques such as Neural Spline Flows (NSFs) and Mixture Density Networks (MDNs), NPE enhanced both the accuracy and scalability of simulation-based inference. In comparison to traditional approaches like Approximate Bayesian Computation (ABC) or Sequential Monte-Carlo (SMC)-ABC (Sisson et al., 2018) that rely on repeated simulations and rejection-based schemes, NPE and its derivatives such as Robust NPE (Ward et al., 2022) are more efficient and flexible.
Sequential Neural Posterior Estimation (SNPE) improves NPE’s efficiency by employing sequential training, where initial parameters are sampled from the prior, and the model is iteratively refined by focusing on regions of the parameter space that are most relevant to the parameter inference task. This adaptive sampling allows for more efficient exploration of the posterior distribution. Early versions like SNPE-A (Papamakarios and Murray, 2016) and SNPE-B (Lueckmann et al., 2017) required correction steps to account for changes in the sampling distribution, adding complexity to the process. In contrast, Automatic Posterior Transformation (APT) or SNPE-C (Greenberg et al., 2019), eliminated this need, offering a more robust and streamlined approach by leveraging normalizing flows to refine the posterior iteratively. However, a limitation of sequential methods like APT is that they necessitate multiple rounds of training, extending the total training time.
The recently proposed Jointly Amortized Neural Approximation (JANA) framework (Radev et al., 2023a) addresses posterior and likelihood estimation through joint amortization. JANA employs conditional invertible neural networks (cINNs) for both the posterior and likelihood networks, allowing efficient transformations between the parameter space and latent variables. Additionally, JANA includes a trainable summary network sub-module optimized to extract maximally informative data representations in an end-to-end manner. Unlike previous approaches, JANA follows a fully amortized strategy, allowing the evaluation of normalized densities and generation of conditional random draws for parameter estimation and surrogate modeling. Simformer (Gloeckler et al., ) is a recent approach that leverages a probabilistic diffusion model with transformer architectures to accurately learn complex, high-dimensional posteriors, making it highly flexible. However, the transformer-based architecture is relatively resource-intensive and involves careful optimization.
We compare our proposed approach with four popular SBI methods: JANA (Radev et al., 2023a), GATSBI (Ramesh et al., 2022), APT (Greenberg et al., 2019), and NPE (Papamakarios and Murray, 2016). The motivation is to cover a variety of approaches (e.g., flow-based, adversarial, one-shot, sequential) while noting that our focus is on efficiency and scalability as opposed to solely the posterior approximation accuracy. We refer the reader to (Zammit-Mangion et al., 2024) and (Radev et al., 2023a) for further reading on deep learning methods for SBI.
3 Simulation Based Inference
In the Bayesian framework, the aim is to infer the posterior distribution $p(\boldsymbol{\theta}\mid\mathbf{y})$ over the model parameters $\boldsymbol{\theta}$, given the observed data $\mathbf{y}$. From Bayes’ theorem, the posterior distribution can be expressed as:
$p(\boldsymbol{\theta}\mid\mathbf{y})=\frac{p(\mathbf{y}\mid\boldsymbol{\theta})p(\boldsymbol{\theta})}{p(\mathbf{y})},$ (1)
[…]
We propose two models for amortized inference: a Conditional Prior VAE (CP-VAE) and an Unconditional Prior VAE (UP-VAE). Both models utilize latent variables to capture hidden structures in the data. The Conditional Prior VAE uses a multivariate prior network that conditions on observed data, enabling adaptive inference that improves generalization across posterior queries. The Unconditional Prior VAE, in contrast, employs a simpler Gaussian prior with unit variance, while still capturing complex posterior distributions. Leveraging the variational inference framework, both models achieve competitive performance across a range of tasks, including bimodal and high-dimensional settings.
3.1 Conditional Prior VAE (CP-VAE)
The Conditional Prior Variational Autoencoder (CP-VAE) is an extension of the standard Conditional Variational Autoencoder (CVAE) (Sohn et al., 2015; Ivanov et al., 2019) specifically designed for simulation-based inference. We begin by introducing an auxiliary latent variable $\mathbf{z}$ to capture complex structures and dependencies within the joint distribution $p(\boldsymbol{\theta},\mathbf{y})$ of parameters $\boldsymbol{\theta}$ and simulated data $\mathbf{y}$. The latent variable $\mathbf{z}$ serves as a learned, low-dimensional summary statistic, efficiently encoding the most salient features of high-dimensional data relevant to characterize the posterior distribution. We formulate a latent variable model $p(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})$ and approximate it using a variational distribution $q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})$, parameterized by $\psi\in\Phi$, where $\Phi$ represents a family of distributions. The goal is to minimize the Kullback-Leibler (KL) divergence between this variational distribution and the true conditional distribution (Jordan et al., 1999; Kingma and Welling, 2014):
$\min_{\psi\in\Phi}D_{KL}\bigg{(}q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})\|p(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})\bigg{)}.$ (2)
Using Bayes’ theorem and the chain rule of probability, the conditional distribution $p(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})$ becomes:
$p(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})=\frac{p(\boldsymbol{\theta}\mid\mathbf{z},\mathbf{y})p(\mathbf{z}\mid\mathbf{y})}{p(\boldsymbol{\theta}\mid\mathbf{y})}.$ (3)
Substituting (3) into (2), and denoting the KL divergence $D_{KL}(q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})\|p(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta}))$ as $\mathcal{D}$, we obtain:
$\mathcal{D}$ $=\mathbb{E}_{\mathbf{z}\sim q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})}\bigg{[}\log q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})-\log p(\mathbf{z}\mid\mathbf{y})$ (4)
$\qquad-\log p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})+\log p(\boldsymbol{\theta}\mid\mathbf{y})\bigg{]}.$
Rearranging the components of equation (4), we isolate $\log p(\boldsymbol{\theta}\mid\mathbf{y})$ to derive:
$\log p(\boldsymbol{\theta}\mid\mathbf{y})\geq$ $-D_{KL}\bigg{(}q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})\parallel p(\mathbf{z}\mid\mathbf{y})\bigg{)}$ (5)
$+\mathbb{E}_{\mathbf{z}\sim q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})}\bigg{[}\log p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})\bigg{]}.$
This inequality stems from the fundamental non-negativity property of the KL divergence, establishing a lower bound on the log-posterior $p(\boldsymbol{\theta}\mid\mathbf{y})$. This bound forms the core of the optimization objective in the CP-VAE framework. Maximizing the log-posterior can be conveniently reformulated as minimizing its negative, which is a standard approach in optimization, leading to the following formulation:
$\mathcal{L}(\boldsymbol{\theta},\mathbf{y};\psi)=D_{KL}\bigg{(}q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})\,\|\,p(\mathbf{z}\mid\mathbf{y})\bigg{)}$ (6)
$-\mathbb{E}_{\mathbf{z}\sim q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})}\bigg{[}\log p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})\bigg{]}.$
Figure 1: CP-VAE Architecture. The encoder $q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})$ computes the parameters of the variational distribution for latent variables $\mathbf{z}$, conditioned on paired simulated data $\mathbf{y}$ and parameters $\boldsymbol{\theta}$. The prior network $p(\mathbf{z}\mid\mathbf{y})$ defines a data-dependent prior over $\mathbf{z}$, enforcing latent space regularization via the KL divergence. The decoder $p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})$ generates the distribution parameters for $\boldsymbol{\theta}$ based on $\mathbf{y}$ and $\mathbf{z}$. Both $\mathbf{z}$ and $\boldsymbol{\theta}$ are sampled using the reparameterization trick, ensuring end-to-end differentiability.
The key advantage of this formulation lies in the prior distribution $p(\mathbf{z}\mid\mathbf{y})$ being conditioned on the observed data $\mathbf{y}$, allowing the model to adaptively shape its latent representation based on observed data, a crucial feature for effective amortized inference. This adaptive prior serves as a powerful regularization mechanism, ensuring that the latent space accurately captures the relevant features from the joint distribution $p(\boldsymbol{\theta},\mathbf{y})$. As a result, this approach delivers more precise posterior approximations, especially when dealing with data that exhibit complex conditional dependencies, which a fixed prior cannot adequately capture. Such adaptability is crucial for models aiming to generalize across a wide range of queries, making the CP-VAE framework well-suited for real-world applications with intricate data dependencies (Sohn et al., 2015).
The loss function presented in (6) is generally intractable due to the complexity of computing both the KL divergence between the variational distribution $q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})$ and the true conditional prior $p(\mathbf{z}\mid\mathbf{y})$, as well as the expectation over the latent variable $\mathbf{z}$. To make the optimization process tractable, we approximate the prior distribution $p(\mathbf{z}\mid\mathbf{y})$, the encoder distribution $q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})$, and the decoder distribution $p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})$ using multivariate Gaussian distributions, each characterized by a mean vector and a covariance matrix. Since each dimension is assumed to be independent, we use a diagonal covariance matrix, which simplifies the calculation of KL divergence. The architecture of the model is illustrated in Figure 1. Detailed derivation of the loss function, along with further explanations of the underlying assumptions, are provided in Section 1 of the Appendix.
To perform amortized inference with CP-VAE for a new observation $\mathbf{y}_{0}$, we first sample $N$ latent variables $\{\mathbf{z}^{(i)}\}_{i=1}^{N}$ from the conditional prior $p(\mathbf{z}\mid\mathbf{y}_{0})$ using the reparameterization trick (Kingma and Welling, 2014). For each latent variable $\mathbf{z}^{(i)}$, the decoder generates the mean and variance of the parameters, from which $\boldsymbol{\theta}^{(i)}$ is then sampled via reparameterization. By aggregating a large number of such samples, we can empirically approximate $p(\boldsymbol{\theta}\mid\mathbf{y}_{0})$, enabling statistical analysis and inference of $\boldsymbol{\theta}$ based on the observed data. The conditional prior, which is dependent on $\mathbf{y}_{0}$, allows the model to adapt its latent representation to the specific data, while the generative decoder maps latent variables to the parameter space, enabling CP-VAE to model complex, data-dependent relationships.
Figure 2: UP-VAE architecture. The encoder $q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})$ outputs the mean and variance of the approximate distribution, from which $\mathbf{z}$ is sampled (reparameterization trick). The KL divergence is computed between the encoder output and a standard normal prior. The theta decoder $p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})$ outputs the parameters for $\boldsymbol{\theta}$, and the data decoder $p(\mathbf{y}\mid\mathbf{z})$ outputs the parameters for $\mathbf{y}$ (reparameterization trick).
3.2 Unconditional Prior VAE (UP-VAE)
A simpler alternative optimization scheme is also proposed, that employs an unconditional prior, in contrast to the earlier approach where the prior was conditioned on the observed data $\mathbf{y}$. Conditioning the prior on $\mathbf{y}$ is beneficial for amortized inference, but adds additional complexity to the model as the prior now depends on the data, potentially increasing the computational burden and complicating inference. Moreover, a data-dependent prior can become overly tailored to the training data, reducing the model’s ability to generalize to new observations. By utilizing an unconditional prior, we simplify the model architecture and promote better generalization by preventing overfitting to training data. To address this, we use Bayes’ theorem to reformulate the conditional distribution $p(\mathbf{z}\mid\boldsymbol{\theta},\mathbf{y})$ as:
$p(\mathbf{z}\mid\boldsymbol{\theta},\mathbf{y})=\frac{p(\boldsymbol{\theta},\mathbf{y}\mid\mathbf{z})p(\mathbf{z})}{p(\boldsymbol{\theta},\mathbf{y})}.$ (7)
Substituting (7) into the KL divergence formulation in (2), and denoting $D_{KL}(q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})\|p(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta}))$ as $\mathcal{D}$,
$\mathcal{D}$ $=\mathbb{E}_{\mathbf{z}\sim q_{\psi}}\bigg{[}\log q_{\psi}(\mathbf{z}\mid\boldsymbol{\theta},\mathbf{y})-\log p(\mathbf{z})\bigg{]}$ (8)
$\quad-\mathbb{E}_{\mathbf{z}\sim q_{\psi}}\bigg{[}\log p(\boldsymbol{\theta},\mathbf{y}\mid\mathbf{z})-\log p(\boldsymbol{\theta},\mathbf{y})\bigg{]}.$
Applying the chain rule to the joint probability $p(\boldsymbol{\theta},\mathbf{y})=p(\mathbf{y})\,p(\boldsymbol{\theta}\mid\mathbf{y})$ and recognizing that $\log p(\mathbf{y})$ is a constant with respect to $\boldsymbol{\theta}$, along with the non-negativity property of the KL divergence, we obtain:
$\log p(\boldsymbol{\theta}\mid\mathbf{y})$ $\geq\mathbb{E}_{\mathbf{z}\sim q_{\psi}}\bigg{[}\log p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})+\log p(\mathbf{y}\mid\mathbf{z})\bigg{]}$ (9)
$\quad-D_{KL}\bigg{(}q_{\psi}(\mathbf{z}\mid\boldsymbol{\theta},\mathbf{y})\,\|\,p(\mathbf{z})\bigg{)}.$
Similar to (5), minimizing the negative of the right-hand side of the derived inequality is equivalent to minimizing the negative log-posterior. This aligns with our primary objective and thus, the final optimization problem can be expressed as:
$\mathcal{L}(\boldsymbol{\theta},\mathbf{y};\psi)=$ $\,D_{KL}\big{(}q_{\psi}(\mathbf{z}\mid\boldsymbol{\theta},\mathbf{y})\,\|\,p(\mathbf{z})\big{)}$ (10)
$-\mathbb{E}_{\mathbf{z}\sim q_{\psi}}\big{[}\log p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})\big{]}$
$-\mathbb{E}_{\mathbf{z}\sim q_{\psi}}\big{[}\log p(\mathbf{y}\mid\mathbf{z})\big{]}.$
This formulation employs two decoder networks: one for reconstructing the parameters $\boldsymbol{\theta}$, and another for reconstructing the observed data $\mathbf{y}$. The data decoder introduces an auxiliary loss, which complements the primary loss from the parameter decoder. This auxiliary objective helps the model capture latent space features that are informative for both $\boldsymbol{\theta}$ and $\mathbf{y}$ (Sheth and Kahou, 2024). However, for amortized inference, only the parameter decoder is needed, as it directly generates the parameter estimates conditioned on the observed data. Similar to CP-VAE, we approximate the prior $p(\mathbf{z})$, the encoder $q_{\psi}(\mathbf{z}\mid\boldsymbol{\theta},\mathbf{y})$, the parameter decoder $p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})$, and the data decoder $p(\mathbf{y}\mid\mathbf{z})$ using Gaussian distributions, simplifying the computation of intractable terms. The architecture of the model is depicted in Figure 2. The derivation of the loss function and the model’s distributional assumptions are provided in Section 2 of the Appendix.
Amortized inference using UP-VAE proceeds by first sampling $N$ latent variables $\{\mathbf{z}^{(i)}\}_{i=1}^{N}$ from the unconditional prior $p(\mathbf{z})$. These latent variables, along with the observed data $\mathbf{y}_{0}$, are then input into the trained theta decoder $p(\boldsymbol{\theta}\mid\mathbf{y}_{0},\mathbf{z}^{(i)})$, which outputs the parameters of the theta distribution from which $\boldsymbol{\theta}^{(i)}$ are sampled. Aggregating these samples empirically approximates $p(\boldsymbol{\theta}\mid\mathbf{y}_{0})$ in an efficient manner.
4 Experiments
[…]
4.1 The Two Moons Model
Figure 3: Two moons: true v/s estimated posterior (sim. budget: 10,000). (a) CP-VAE and (b) UP-VAE
The Two Moons model is a popular benchmark commonly used to evaluate the ability of SBI algorithms to recover multimodal and unconventional posterior distributions (Greenberg et al., 2019; Lueckmann et al., 2021). The challenge herein is to recover the bimodal nature of the posterior, and to capture its distinctive crescent shape. The parameter vector $\boldsymbol{\theta}\sim\mathcal{U}(-1,1)^{2}$ governs the model, and the observed data $\mathbf{x}$ is generated as:
$\mathbf{x}|\boldsymbol{\theta}=\begin{pmatrix}r\cos(a)+0.25\\ r\sin(a)\end{pmatrix}+\frac{1}{\sqrt{2}}\begin{pmatrix}-|\theta_{1}+\theta_{2}|\\ -(\theta_{1}-\theta_{2})\end{pmatrix},$ (11)
for radius $r\sim\mathcal{N}(0.1,0.01^{2})$ and angle $a\sim\mathcal{U}\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$.
Figure 3 shows the recovered posterior using (a) CP-VAE and (b) UP-VAE model formulations. Both models generate samples that align closely with the true posterior, with UP-VAE exhibiting slightly better coverage of the crescent shape. Certain discrepancies, such as outliers, may be attributed to the variational lower bound in the loss function. Tables 1 and 2 present a comparative analysis using the C2ST accuracy and MMD metrics respectively. JANA and APT perform well, accurately characterizing the bimodal posterior. CP-VAE and UP-VAE closely follow, maintaining accuracy with low variance. Notably, both CP-VAE and UP-VAE are training efficient (Table 3), while demonstrating effectiveness in handling complex, unconventional posterior distributions.
Figure 4: SIR model: true v/s estimated posterior distribution. (a) CP-VAE and (b) UP-VAE.
4.2 The Bernoulli Generalized Linear Model
The Bernoulli Generalized Linear Model (GLM) simulates neuronal spiking behavior in response to a time-varying stimulus, driven by a set of covariates (Gonzalves et al., 2020). With a discrete time domain of $T=100$ bins, the model produces observed data in the form $\mathbf{y}=[y_{i}]_{i=1}^{10}$. Each component $y_{i}$ aggregates spiking activity across time intervals, capturing both overall spiking frequency and spatial patterns of neuronal activity. The model is parameterized by a vector $\boldsymbol{\theta}\in\mathbb{R}^{10}$, consisting of a scalar bias term $\beta$ and an auxiliary vector $\mathbf{f}\in\mathbb{R}^{9}$:
$\boldsymbol{\theta}=\begin{bmatrix}\beta\\ \mathbf{f}\end{bmatrix},\quad\beta\sim\mathcal{N}(0,2),\quad\mathbf{f}\sim\mathcal{N}\left(\mathbf{0},\left(\mathbf{F}^{\top}\mathbf{F}\right)^{-1}\right),$ (12)
where the matrix $\mathbf{F}$ enforces smoothness by penalizing second-order differences in $\mathbf{f}$, defined as:
$F_{ij}=\begin{cases}1,&j=i-2,\\ -2,&j=i-1,\\ 1+\sqrt{\frac{i-1}{9}},&j=i,\\ 0,&\text{otherwise.}\end{cases},\quad 1\leq i\leq 9,$ (13)
For each time bin $i$, the binary spiking variable $z_{i}$ is sampled from a Bernoulli distribution, where the spiking probability is modeled using a logistic function applied to the linear combination of covariates $\mathbf{v}_{i}$ and the parameter vector, expressed as:
$z_{i}\sim\text{Bernoulli}\left(\sigma(\mathbf{v}_{i}^{\top}\mathbf{f}+\beta)\right),\quad\sigma(t)=\frac{1}{1+e^{-t}}.$ (14)
Here, $\mathbf{v}_{i}\in\mathbb{R}^{9}$ represents white noise inputs (covariates) drawn from $\mathcal{N}(0,\mathbf{I})$. The design matrix $\mathbf{V}=\left[\mathbf{v}_{i}\right]_{i=1}^{T}$ aggregates these inputs over time bins $T$. The model computes summary statistics $\mathbf{y}\in\mathbb{R}^{10}$, where the total number of spikes is represented by $y_{1}=\sum_{i=1}^{T}z_{i}$, and the remaining components, $y_{2},\dots,y_{10}$, are spike-triggered averages of the covariates, normalized by $y_{1}$, defined as $y_{2:10}=\frac{1}{y_{1}}\mathbf{V}\mathbf{z}$. The summary statistics provide a compact representation of the simulated spiking activity, capturing the overall frequency and the influence of covariates on the spiking process.
Figure 5 presents the posterior distributions of $\theta_{1}$ through $\theta_{10}$ for CP-VAE and UP-VAE. In both cases, the posterior distributions align well with the true posteriors, though CP-VAE exhibits a better fit as compared to UP-VAE. CP-VAE achieves the lowest C2ST accuracy (Table 1) and MMD values (Table 2), indicating a closer match to the true posterior. UP-VAE follows closely behind on the C2ST metric, while NPE is the second-best method on the MMD metric. Both CP-VAE and UP-VAE are substantially faster to train (Table 3) showcasing efficiency and scalability for this test problem involving 10 parameters.
4.3 The SIR Model
The Susceptible-Infected-Recovered (SIR) model is a widely used mathematical framework in epidemiology for modeling the spread of infectious diseases within a population. This model tracks the evolution of the number of individuals in three distinct compartments: susceptible ($S$), infected ($I$), and recovered ($R$). The dynamics of disease spread are governed by a non-linear dynamical system characterized by two key parameters: the contact rate ($\beta$) and the recovery rate ($\gamma$). The model is described by the following set of ordinary differential equations:
$\frac{d}{dt}\begin{pmatrix}S\\ I\\ R\end{pmatrix}=\begin{pmatrix}-\beta S\\ \beta S-\gamma I\\ \gamma I\end{pmatrix}\frac{I}{N},$ (15)
where $N$ is the total population size. The parameters are sampled as: $\beta\sim\text{LogNormal}(\log(0.4),0.5)$ and $\gamma\sim\text{LogNormal}(\log(1/8),0.2)$.
Figure 4 presents the posterior distributions estimated using CP-VAE and UP-VAE, respectively. Both models effectively capture complex, data-dependent relationships, resulting in accurate posterior approximations. Both CP-VAE and UP-VAE performed well on the C2ST (Table 1) and MMD metrics (Table 2), while being highly efficient (training times in Table 3).
Figure 5: Bernoulli GLM: true v/s estimated posterior distribution. (a) CP-VAE and (b) UP-VAE.
Table 1: C2ST metric for different methods across simulation budgets (mean $\pm$ std. dev.).
Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.7770 (± 0.0630) 0.6634 (± 0.0503) 0.7467 (± 0.1036) 0.9999 (± 0.0001) 0.9999 (± 0.0000) 0.9997 (± 0.0005) 0.9990 (± 0.0015) 0.9912 (± 0.0161) 0.9838 (± 0.0157) NPE 0.5749 (± 0.0331) 0.5931 (± 0.0329) 0.5411 (± 0.0195) 0.9121 (± 0.0338) 0.8988 (± 0.0306) 0.8898 (± 0.0565) 0.8378 (± 0.0396) 0.9122 (± 0.0207) 0.9405 (± 0.0125) JANA 0.5206 (± 0.0053) 0.5151 (± 0.0055) 0.5123 (± 0.0032) 0.9678 (± 0.0035) 0.9692 (± 0.0026) 0.9589 (± 0.0034) 0.8697 (± 0.0149) 0.9175 (± 0.0059) 0.9343 (± 0.0055) APT 0.5189 (± 0.0080) 0.5202 (± 0.0114) 0.5147 (± 0.0125) 0.8404 (± 0.0107) 0.8328 (± 0.0156) 0.8353 (± 0.0164) 0.9433 (± 0.0076) 0.9434 (± 0.0061) 0.9457 (± 0.0068) CP-VAE 0.6443 (± 0.0288) 0.6072 (± 0.0108) 0.6346 (± 0.0303) 0.6955 (± 0.0239) 0.6697 (± 0.0361) 0.6427 (± 0.0177) 0.8524 (± 0.0719) 0.8976 (± 0.0121) 0.8857 (± 0.0330) UP-VAE 0.6510 (± 0.0310) 0.6565 (± 0.0364) 0.6212 (± 0.0252) 0.8270 (± 0.0249) 0.8311 (± 0.0216) 0.7819 (± 0.0692) 0.8390 (± 0.0378) 0.8731 (± 0.0482) 0.8893 (± 0.0317)
Table 2: MMD metric for different methods across simulation budgets (mean $\pm$ std. dev.).
Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.1213 (± 0.0650) 0.0582 (± 0.0122) 0.1432 (± 0.1161) 3.6381 (± 0.5428) 4.8299 (± 0.6825) 3.9965 (± 1.8510) 2.9470 (± 0.4891) 2.7502 (± 0.7268) 2.1466 (± 0.5625) NPE 0.0254 (± 0.0025) 0.0241 (± 0.0099) 0.0125 (± 0.0045) 1.0371 (± 0.1830) 0.9379 (± 0.1001) 0.8675 (± 0.1945) 0.6611 (± 0.2236) 1.0312 (± 0.1255) 1.1903 (± 0.1007) JANA 0.0060 (± 0.0013) 0.0068 (± 0.0017) 0.0047 (± 0.0009) 3.3232 (± 0.0932) 3.1488 (± 0.0280) 2.9010 (± 0.0878) 0.8495 (± 0.1254) 1.1077 (± 0.0336) 1.1421 (± 0.0392) APT 0.0073 (± 0.0030) 0.0069 (± 0.0030) 0.0037 (± 0.0012) 0.7383 (± 0.0853) 0.4944 (± 0.0663) 0.5808 (± 0.0539) 1.1003 (± 0.0712) 1.1093 (± 0.0311) 1.1291 (± 0.0481) CP-VAE 0.0382 (± 0.0110) 0.0369 (± 0.0088) 0.0276 (± 0.0050) 0.1966 (± 0.1375) 0.1557 (± 0.0663) 0.0942 (± 0.0547) 0.8792 (± 0.3327) 1.0158 (± 0.0566) 0.9990 (± 0.1923) UP-VAE 0.0736 (± 0.0097) 0.0795 (± 0.0151) 0.0544 (± 0.0150) 1.5420 (± 0.4930) 1.4748 (± 0.2623) 1.1160 (± 0.7496) 0.5168 (± 0.1193) 0.6884 (± 0.2044) 0.6487 (± 0.1087)
5 Discussion
Table 3: Training time for different methods across simulation budgets (mean $\pm$ std. dev. in minutes).
Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 117.67 (± 2.43) 164.90 (± 57.10) 249.50 (± 105.05) 118.62 (± 0.18) 179.69 (± 49.96) 253.48 (± 113.38) 227.16 (± 1.44) 241.24 (± 7.58) 250.47 (± 2.22) NPE 7.04 (± 0.21) 14.21 (± 0.04) 20.95 (± 0.37) 12.17 (± 0.17) 24.96 (± 0.64) 37.43 (± 0.22) 20.10 (± 5.45) 29.62 (± 3.74) 44.92 (± 9.24) JANA 5.66 (± 0.10) 9.48 (± 0.12) 14.34 (± 0.11) 7.14 (± 0.05) 12.23 (± 0.07) 19.00 (± 0.19) 22.98 (± 0.32) 39.17 (± 2.48) 61.89 (± 5.41) APT 90.31 (± 8.99) 186.17 (± 13.79) 268.28 (± 32.88) 101.39 (± 7.89) 197.30 (± 9.32) 267.38 (± 21.93) 40.15 (± 3.90) 73.74 (± 6.25) 112.51 (± 2.51) CP-VAE 2.23 (± 0.81) 3.75 (± 0.80) 4.99 (± 0.83) 1.76 (± 0.27) 3.64 (± 0.14) 4.93 (± 0.47) 2.77 (± 0.40) 6.06 (± 1.09) 8.68 (± 0.83) UP-VAE 1.98 (± 0.33) 3.27 (± 0.58) 5.35 (± 1.82) 2.10 (± 0.50) 4.43 (± 0.61) 6.57 (± 1.04) 4.44 (± 0.88) 6.71 (± 1.97) 9.60 (± 1.66)
The results are now evaluated with respect to the following important considerations in SBI.
Posterior estimation accuracy: JANA and APT hold an advantage, as JANA is able to capture complex dependencies between parameters, owing to the autoregressive factorization of the joint posterior as a series of conditional distributions; while APT leverages sequential posterior refinement to enhance accuracy.
Computational efficiency: Adversarial training presents well known challenges, which reflect in relatively high training times for GATSBI. The sequential nature of APT adds computational expense over NPE and CP/UP-VAE. Autoregressive models can be computationally expensive to train and evaluate as estimating each parameter depends on previously estimated parameters – however, JANA is able to closely match NPE’s training times for chosen hyperparameters. CP-VAE and UP-VAE are inherently simple, and do not require external summary statistics, aiding scalability and efficient training (Table 3).
Robustness/stability: GATSBI suffers from the instabilities associated with the adversarial setting (e.g., high MMD std. dev. in Table 2). JANA and APT proved to be very robust, with low std. dev. across test problems, followed by CP/UP-VAE and NPE.
Flexibility/approximating complex posteriors: APT holds an advantage over NPE due to its iterative refinement. JANA is particularly suited for posteriors with dependencies between parameters. However, the flexibility of normalizing flow based approaches (NPE, APT, JANA) is hindered by the invertibility requirement. GATSBI is suitable for complex, high-dimensional problems where other approaches may struggle and learning the posterior implicitly becomes important. Both CP-VAE and UP-VAE are highly flexible and adaptable with respect to the architecture.
CP-VAE v/s UP-VAE: CP-VAE’s conditional prior $p(\mathbf{z}\mid\mathbf{y})$ adapts to observed data, allowing flexibility and control for improved posterior accuracy. This is of particular interest in high-dimensional problems. UP-VAE’s simple Gaussian prior $p(\mathbf{z})$ allows for efficient latent space sampling, easier regularization, and a consistent latent space, leading to more stable training, making it efficient and effective for simpler problems.
The proposed CP-VAE and UP-VAE models complement existing approaches by virtue of approximating the posterior distribution in a computation-efficient manner (as evidenced in Table 3). This makes the models well-suited for challenging problems where computational efficiency is of key consideration (e.g., large number of parameters to infer, rapid model exploration, etc.). While UP-VAE uses a simpler prior, its data decoder is computationally more expensive than the conditional prior in CP-VAE, which reflects in faster training times for CP-VAE for most problems (Table 3). While we note that the training times in Table 3) can be affected by several factors including implementation and hyperparameter optimization, the comparison provides an estimate of relative computational complexity on the same hardware for reference implementations (taken from the original papers).
6 Conclusion
This paper introduces two latent variable models, CP-VAE and UP-VAE, for amortized inference in likelihood-free simulation-based inference, with a view on efficiency, scalability, interpretability and robustness. The proposed models are validated against state-of-the-art normalizing flow and adversarial training-based approaches on benchmark test problems. Both CP-VAE and UP-VAE deliver comparable performance to the state-of-the-art while being highly computation efficient and scalable. Future enhancements could involve integrating more sophisticated architectures (e.g., Convolutional Neural Networks) to better capture spatial and temporal dependencies within the data. Exploring diverse prior distributions, such as truncated Gaussian priors, may further improve the models’ ability to approximate complex posteriors.
Acknowledgements
The computations/data handling were enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre and by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Chalmers e-Commons at Chalmers, and Uppsala Multidisciplinary Center for Advanced Computational Science (UPPMAX) at Uppsala University, partially funded by the Swedish Research Council through grant agreement nos. 2022-06725 and 2018-05973. AH and PS acknowledge support from the Swedish Research Council through grant agreement nos. 2023-05167 and 2023-05593 respectively.